# Quasi-symmetry Constrained Spin Ferromagnetism in Altermagnets

### v1

@@ Abstract
where the dispersion is $E_{\alpha=\pm,\beta=\pm}=\alpha\Big{(}N_{x}^{2}+\tilde{E}^{2}_{\bf k}+\beta 2N_{x}\sqrt{t_{z,{\bf k}}^{2}+\lambda_{y,{\bf k}}^{2}+\lambda_{z,{\bf k}}^{2}}\Big{)}^{1/2}$. This term reveals that it is the SOC component parallel to the Néel vector that determines the AHE. This relationship between the SOC direction and non-vanishing AHE can also be understood using SOC-based quasi-symmetry arguments like those given above. Specifically, the AHE measures current and voltage response, and hence is even under spin-rotational symmetries. For the Néel order along $\hat{x}$, when only the $\lambda_{y}$ ($\lambda_{z}$) SOC is kept, the Néel order is odd under the resulting quasi-spin-rotational symmetry $C_{2y}^{\rm spin}$ ($C_{2z}^{\rm spin}$). This prevents a $\lambda_{y}$ ($\lambda_{z}$)-linear SOC contribution to the AHE. When only the $\lambda_{x}$ SOC is kept, the Néel order is even under the resulting $C_{2x}^{\rm spin}$ symmetry, and therefore the $\lambda_{x}$-linear contribution to AHE is allowed by the quasi-symmetry. It naturally follows that the AHE is given by the SOC component that is parallel to the Néel vector. We note that this quasi-symmetry argument also applies to ferromagnetic order and this is confirmed by explicit microscopic calculations in the SM Roig et al. .
@@
with the function $L({\bf k})$ defined in Eq. (8). The sign of these coefficients fixes the easy axis for a specific AM material. For instance, focusing on SG 136, $\lambda_{z,{\bf k}}$ can be ignored as it is smaller than $\lambda_{x,{\bf k}}$ and $\lambda_{y,{\bf k}}$ Roig et al. (2024). Hence, in general we expect that $s_{2}>0$ and, as a consequence, the easy axis is out of plane. However, Eqs. (11)-(12) also reveal that the moment direction may switch depending on the Fermi energy. In particular, when the Lindhard function (interband term in $L({\bf k})$) dominates over the density of states (intraband term), $L({\bf k})$ can change sign leading to $s_{2}<0$ and in-plane moment orientation. The switch of the AM moments as a function of the Fermi energy from the $c$-axis to the in-plane direction has been reported in Ref. Feng et al. (2022) for RuO2. As shown in the SM section, calculations of $s_{1}$ and $s_{2}$ for a minimal model of RuO2 indeed yield $s_{1}=0$, $s_{2}>0$. In addition, application to FeSb2 reveal that $s_{1}>0$, $s_{2}<0$, i.e. moments aligned along the in-plane $y$-axis, in agreement with DFT studies Mazin et al. (2021). A systematic application of the magnetic anisotropy energy based on Eqs. (11)-(12) to other AM is beyond the scope of this work, and constitute an interesting future project.
Conclusions.– In summary we have applied recently-developed realistic microscopic models for AM to derive relevant Landau coefficients of the free energy, focusing on the coupling between magnetization and Néel order, and magnetic anisotropy energies. The results explain an observed strong material-dependence of the SOC-induced weak FM. We stress that the weak FM described in this work refers to the spin moment. For materials where this moment is forbidden to linear order, e.g. the rutiles, secondary orders become relevant, and the orbital magnetic moment may also yield contributions to the small but finite net magnetization Šmejkal et al. (2020); Jo et al. (2024). Finally, we discovered a general quasi-symmetry enabled by the SOC that allows to determine for which AM symmetries the induced FM spin moment is large or vanishingly small.
@@ Acknowledgements.
Acknowledgements.– M. R. acknowledges support from the Novo Nordisk Foundation grant NNF20OC0060019. A. K. acknowledges support by the Danish National Committee for Research Infrastructure (NUFI) through the ESS-Lighthouse Q-MAT. D. F. A. and Y. Y. were supported by the National Science Foundation Grant No. DMREF 2323857.

### v2

@@ Abstract
The relationship between the SOC direction and non-vanishing AHE in Eq. (3) can also be understood using SOC-based quasi-symmetry arguments like those given above, which do not depend on the form of the microscopic model. Specifically, the AHE measures current and voltage response, and hence is even under spin-rotational symmetries. For the Néel order along $\hat{x}$, when only the $\lambda_{y}$ ($\lambda_{z}$) SOC is kept, the Néel order is odd under the resulting quasi-spin-rotational symmetry $[C_{2y}||E]$ ($[C_{2z}||E]$). This prevents a $\lambda_{y}$ ($\lambda_{z}$)-linear SOC contribution to the AHE. When only the $\lambda_{x}$ SOC is kept, the Néel order is even under the resulting $[C_{2x}||E]$ symmetry, and therefore the $\lambda_{x}$-linear contribution to AHE is allowed by the quasi-symmetry. It naturally follows that the AHE is given by the SOC component that is parallel to the Néel vector.
@@
with the function $L({\bf k})$ defined in Eq. (9). The sign of these coefficients fixes the easy axis for a specific AM material. For instance, focusing on SG 136, $\lambda_{z,{\bf k}}$ can be ignored as it is smaller than $\lambda_{x,{\bf k}}$ and $\lambda_{y,{\bf k}}$ Roig et al. (2024). Hence, in general we expect that $s_{2}>0$ and, as a consequence, the easy axis is out of plane. However, Eqs. (11)-(12) also reveal that the moment direction may switch depending on the Fermi energy. In particular, when the Lindhard function (interband term in $L({\bf k})$) dominates over the density of states (intraband term), $L({\bf k})$ can change sign leading to $s_{2}<0$ and in-plane moment orientation. The switch of the AM moments as a function of the Fermi energy from the $c$-axis to the in-plane direction has been reported in Ref. Feng et al. (2022) for RuO2. As shown in the SM, calculations of $s_{1}$ and $s_{2}$ for a minimal model of RuO2 indeed yield $s_{1}=0$, $s_{2}>0$. In addition, application to FeSb2 reveal that $s_{1}>0$, $s_{2}<0$, i.e. moments aligned along the in-plane $y$-axis, in agreement with DFT studies Mazin et al. (2021). A systematic application of the magnetic anisotropy energy based on Eqs. (11)-(12) to other AM is beyond the scope of this work, and constitute an interesting future project.
Conclusions.– In summary, we have applied recently-developed realistic microscopic models for AM to derive relevant Landau coefficients of the free energy, focusing on the coupling between magnetization and Néel order, and magnetic anisotropy energies. The results explain the generic large AHE and the observed strong material-dependence of the SOC-induced weak FM. We stress that the weak FM described in this work refers to the spin moment. For materials where this moment is forbidden to linear order, e.g. the rutiles, secondary orders become relevant, and the orbital magnetic moment may also yield contributions to the small but finite net magnetization Šmejkal et al. (2020); Jo et al. (2024). Finally, we discovered a general quasi-symmetry enabled by the SOC that is model-independent and allows to determine for which AM symmetries the induced FM spin moment is large or vanishingly small.
@@ Acknowledgements.
Acknowledgements.– M. R. acknowledges support from the Novo Nordisk Foundation grant NNF20OC0060019. A. K. acknowledges support by the Danish National Committee for Research Infrastructure (NUFI) through the ESS-Lighthouse Q-MAT. D. F. A. and Y. Y. were supported by the National Science Foundation Grant No. DMREF 2323857. Work at UWM was also supported by a grant from the Simons Foundation (SFI-MPS-NFS-00006741-02, D.F.A and M.R.).
Appendix A End Matter
Appendix: General SOC-enabled quasi-symmetry– As introduced in the main text, the quasi-symmetry generated by SOC when two of the SOC components vanish allows us to determine which Landau coefficients are linear in one of $\lambda_{x,y,z}$. As illustrated in Fig. 3, the resulting symmetry group, denoted here as the uniaxial spin space group, has lower symmetry than the spin space group, but higher symmetry than the magnetic space group. Here, we develop a more general quasi-symmetry criterion that addresses the following question: can a Landau coefficient or response function host a contribution from $\lambda_{x}^{n_{x}}\lambda_{y}^{n_{y}}\lambda_{z}^{n_{z}}$, for given integers $(n_{x},n_{y},n_{z})\geq 0$?
We would like to consider a general Hamiltonian that can extend beyond the minimal models discussed in this paper. SOC terms can always be grouped into terms proportional to spin operators $\sigma_{x}$, $\sigma_{y}$, and $\sigma_{z}$. Terms in each group can carry different momentum and orbital dependence, with overall SOC strengths $\lambda_{x}$, $\lambda_{y}$, and $\lambda_{z}$:
$H_{SOC}=$ $\,\ \lambda_{x}(a_{1,{\bf k}}\hat{A_{1}}+a_{2,{\bf k}}\hat{A_{2}}+...)\sigma_{x}$
$+\lambda_{y}(b_{1,{\bf k}}\hat{B_{1}}+b_{2,{\bf k}}\hat{B_{2}}+...)\sigma_{y}$
$+\lambda_{z}(c_{1,{\bf k}}\hat{C_{1}}+c_{2,{\bf k}}\hat{C_{2}}+...)\sigma_{z},$ (13)
with momentum-dependent coefficients $a_{i},b_{i},c_{i}=O(1)$ and $\hat{A_{i}}$, $\hat{B_{i}}$, $\hat{C_{i}}$ some operators acting on sublattice and/or orbital space. The quasi-symmetry argument is based on analyticity of the response functions. If the DOS is not concentrated at singular regions such as band crossings, response functions should be an analytical function of $\lambda_{x,y,z}$. Cross-terms like $\sqrt{\lambda_{x}\lambda_{y}}$ are not allowed. For the minimal model, this property is explicitly shown in the normal-state Green’s function Eq. (6), where the projection operator is a linear combination of the three SOC terms.
Figure 3: Classification scheme of the spin space group and magnetic space group in terms of the strength and powers of SOC, illustrating when the quasi-symmetries emerge. In particular, when two of the SOC components vanish, the resulting symmetry group (which we denote as the uniaxial spin space group) has spin rotational invariance around the SOC direction.
We now interpret the three groups of SOC terms as distinct symmetry-breaking order parameters in a SOC-free system (spin-space group). The spin-space group hosts decoupled orbital crystal symmetries and full spin-rotational symmetries. General order parameters are denoted by $\vec{O}_{R,i}$, where R is the IR under orbital rotations and the index $i=1,2,\dots$ reflects its dimensionality. The vector notation indicates that each $\vec{O}_{R,i}$ also transforms as a spin vector under spin rotations.
Each group of SOC terms corresponds to a different symmetry breaking in this high-symmetry system. For example, consider the $\lambda_{z}$ SOC term in an tetragonal system. It breaks spin-rotational symmetry as the spin vector $\sigma_{z}$, and it breaks orbital symmetry according to the 1D IR $A_{2g}\sim xy(x^{2}-y^{2})$. It is then $\vec{O}_{A_{2g}}=(0,0,\Lambda_{z})$. For $\lambda_{x,y}$ SOC terms, they break spin-rotational symmetry as the spin vector $\sigma_{x,y}$, and they break orbital symmetry according to the 2D IR $E_{g}\sim\{xz,yz\}$. They are then $\vec{O}_{E_{g},1}=(\Lambda_{xy},0,0)$ and $\vec{O}_{E_{g},2}=(0,\Lambda_{xy},0)$.
If the coefficient of a Landau term or response function $X$ hosts a contribution from $\lambda_{x}^{n_{x}}\lambda_{y}^{n_{y}}\lambda_{z}^{n_{z}}$, then the term $\Lambda_{x}^{n_{x}}\Lambda_{y}^{n_{y}}\Lambda_{z}^{n_{z}}X$ must be allowed in the Landau theory for the SOC-free system. We note that for the tetragonal system, the contributions from $\lambda_{x,y}^{n_{xy}}\lambda_{z}^{n_{z}}$ enable the term $\Lambda_{xy}^{n_{xy}}\Lambda_{z}^{n_{z}}X$ in the Landau theory for the SOC-free system (since $\lambda_{x}=\lambda_{y}$ in this case). In the SOC-free system, a valid Landau term should be a spin-scalar, and belong to the trivial IR under orbital rotations. This provides a general quasi-symmetry criterion.
Examples–
  1. 1.
    In the SOC-free tetragonal $D_{4h}$ systems, the SOC order parameters are: $\vec{O}_{E_{g},1}=(\Lambda_{xy},0,0)$, $\vec{O}_{E_{g},2}=(0,\Lambda_{xy},0)$ and $\vec{O}_{A_{2g}}=(0,0,\Lambda_{z})$. Ferromagnetic order is $\vec{M}_{A_{1g}}$, and altermagnetic order is $\vec{N}_{P}$, with altermagnetic symmetry $\Gamma_{N}=P$. To have a SOC-linear coupling between an altermagnet and a ferromagnet, $\vec{O}\cdot(\vec{M}\times\vec{N})$ must be allowed. The altermagnetic symmetry thus has to be $A_{2g}$ or $E_{g}$. For two atoms per nonmagnetic unit cell, only $\Gamma_{N}=A_{2g},B_{1g},B_{2g}$ are allowed Roig et al. (2024).
    1. (a)
      For $\Gamma_{N}=A_{2g}$, the Landau term $\vec{O}_{A_{2g}}\cdot(\vec{M}_{A_{1g}}\times\vec{N}_{A_{2g}})=\Lambda_{z}(M_{x}N_{y}-M_{y}N_{x})$ is allowed. Hence the $\lambda_{z}$-linear contribution to $M_{x}N_{y}-M_{y}N_{x}$ is nonzero.
    2. (b)
      For $\Gamma_{N}=B_{2g}$ and Landau term $M_{x}N_{y}+M_{y}N_{x}$, SOC-linear terms are forbidden. Since $E_{g}\otimes E_{g}\otimes B_{2g}$ contains the trivial IR, the quadratic $\lambda_{xy}^{2}$ contribution is allowed. An example Landau term is $(\vec{O}_{E_{g},1}\times\vec{M})\cdot(\vec{N}\times\vec{O}_{E_{g},2})+(\vec{O}_{E_{g},2}\times\vec{M})\cdot(\vec{N}\times\vec{O}_{E_{g},1})=\Lambda_{xy}^{2}(M_{x}N_{y}+M_{y}N_{x})$.
    3. (c)
      For $\Gamma_{N}=B_{1g}$ and Landau term $M_{x}N_{x}-M_{y}N_{y}$, SOC-linear terms are forbidden. Since $E_{g}\otimes E_{g}\otimes B_{1g}$ contains the trivial IR, the quadratic $\lambda_{xy}^{2}$ contribution is allowed. An example Landau term is $(\vec{O}_{E_{g},1}\times\vec{M})\cdot(\vec{N}\times\vec{O}_{E_{g},1})-(\vec{O}_{E_{g},2}\times\vec{M})\cdot(\vec{N}\times\vec{O}_{E_{g},2})=\Lambda_{xy}^{2}(M_{x}N_{x}-M_{y}N_{y})$.
  2. 2.
    In the SOC-free hexagonal $D_{6h}$ systems, the SOC order parameters are: $\vec{O}_{E_{1g},1}=(\Lambda_{xy},0,0)$, $\vec{O}_{E_{1g},2}=(0,\Lambda_{xy},0)$ and $\vec{O}_{A_{2g}}=(0,0,\Lambda_{z})$. To have a SOC-linear coupling between altermagnet and ferromagnet, $\vec{O}\cdot(\vec{M}\times\vec{N})$ must be allowed. Thus, $\Gamma_{N}=A_{2g}$ or $E_{1g}$. For two atoms per nonmagnetic unit cell, only $A_{2g}$, $B_{1g}$, and $B_{2g}$ are allowed Roig et al. (2024).
    1. (a)
      The discussion for $\Gamma_{N}=A_{2g}$ is the same as $A_{2g}$ in tetragonal systems.
    2. (b)
      For $\Gamma_{N}=B_{1g}$ (similar for $\Gamma_{N}=B_{2g}$), $MN^{3}$ coupling has no SOC-linear contribution since $B_{1g}\otimes B_{1g}\otimes B_{1g}\otimes\{E_{1g},A_{2g}\}$ has no trivial IR. Similarly, SOC-quadratic contribution is also forbidden since $B_{1g}\otimes B_{1g}\otimes B_{1g}\otimes\{E_{1g},A_{2g}\}\otimes\{E_{1g},A_{2g}\}$ has no trivial IR. SOC-cubic contribution is allowed as $B_{1g}\otimes B_{1g}\otimes B_{1g}\otimes E_{1g}\otimes E_{1g}\otimes E_{1g}$ has trivial IR. The SOC dependence is cubic in $\lambda_{xy}$.
  3. 3.
    In the SOC-free cubic $O_{h}$ systems, the SOC order parameters are: $\vec{O}_{T_{1g},1}=(\Lambda,0,0)$, $\vec{O}_{T_{1g},2}=(0,\Lambda,0)$ and $\vec{O}_{T_{1g},3}=(0,0,\Lambda)$. For $\Gamma_{N}=A_{2g}$, $MN^{3}$ coupling has no SOC-linear contribution since $A_{2g}\otimes A_{2g}\otimes A_{2g}\otimes T_{1g}$ has no trivial IR. Similarly, a SOC-quadratic contribution is also forbidden since $A_{2g}\otimes A_{2g}\otimes A_{2g}\otimes T_{1g}\otimes T_{1g}$ has no trivial IR. Hence, the leading contribution is cubic in SOC, as $A_{2g}\otimes A_{2g}\otimes A_{2g}\otimes T_{1g}\otimes T_{1g}\otimes T_{1g}$ contains the trivial IR.
[…]